What is the molar mass of an unknown gas with a density of 2.50 g/L at 1.00 atm and 25.0 °C?
Answer #1

P = 1.0atm
T= 25.0 oC
= (25.0+273) K
= 298 K
density = 2.50 g/L
Lets derive the equation to be used
use:
p*V=n*R*T
p*V=(mass/molar mass)*R*T
p*molar mass=(mass/V)*R*T
p*molar mass=density*R*T

Put Values:
1 atm * MM = 2.5g/L * 0.08206 atm.L/mol.K *298.0 K
MM = 61.1347 g/mol
Answer: 61.13 g/mol